OpenStudy (ggguhys):

The fire chief is throwing a Christmas party for the 24 firefighters, medics, and officers in her volunteer fire department. She finds 6 different Disney characters dressed in protective bunker gear and buys an equal number of each. All packages are wrapped the same way and each member of the department selects his or her own gift. The medical officer selects first, followed by the engineer and then the deputy chief. What is the probability that each of the three will select the Pooh Bear character? A 5/506 B 1/506 C 5/576 D 1/576

OpenStudy (mathmate):

24 members in the department, 6 types of gears, so 4 of each kind. Medical officer chooses 4 Pooh Bear character out of 24, with probability 4/24. Engineer chooses remaining 3 out of 23, with probability of 3/23 Deputy chief chooses 2 remaining out of 22, with probability of 2/22 For all of this to happen, we multiply the individual probabilities, so P(all Pooh Bears) = (4/24)*(3/23)*(2/22)=1/506